OpenStudy (anonymous):

PLS HELP MEDAL IS REWARDED Explain what common logarithms are and how they are used to find the value of a logarithm that is not common.

OpenStudy (anonymous):

A common logarithm is a logarithm to the base 10. LOG(x) = a means that 10^a = x, so LOG(10)=1, LOG(100)=2, etc. To find a logarithm of a different base, b, consider Given x, we want y such that b^y = x. Since b = 10^(LOG(b)) b^y = (10^LOG(b)) ^ y = x We know that (10^m)^n = 10 ^ (m*n) so x = 10 ^ (LOG(b)*y) and LOG(x) = LOG(b)*y so y = LOG(x) / LOG(b)

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

u welcome